OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 2x + 5y = 7 3x + 5y = 3

OpenStudy (helder_edwin):

do u mean CHOOSE or FIND?

OpenStudy (anonymous):

'find', sorry

OpenStudy (helder_edwin):

do you know cramer's rule?

OpenStudy (anonymous):

ehh sort of

OpenStudy (anonymous):

Okay, so in this case, the regular coefficient matrix would be this:\[\begin{bmatrix}2&5\\3&5\end{bmatrix}\] But we want the y determinant, so we replace the y coefficients with the solution values, to obtain this:\[\begin{bmatrix}2&7\\3&3\end{bmatrix}\]Then we find the determinant in the normal way.

OpenStudy (anonymous):

still confused

OpenStudy (helder_edwin):

it is easy \[ \large y=\frac{|D(y)|}{|A|} \] where \[ \large A=\begin{pmatrix} 2 & 5\\ 3 & 5 \end{pmatrix}\qquad\text{and}\qquad D(y)=\begin{pmatrix} 2 & 7\\ 3 & 3 \end{pmatrix} \]

OpenStudy (anonymous):

is the answer -15?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks guys!